OpenStudy (anonymous):

Two oppositely and equally charged particles of equivalent mass 'm' a distance 'd' apart. find the time it takes for the two particles to collide.

OpenStudy (anonymous):

U(x) = qV(x) V(x) = kq/x U = kq^2/x KE = 1/2mv^2 v(x) = sqrt( 2kq^2/mx ) dx/dt = v(x) 1/v(x) dx = dt t = Int[0,1/2d] 1/v(x) dx t = sqrt( mx/8kq^2 ) Pretty sure I did it incorrectly, was wondering what other people thought.

OpenStudy (anonymous):

eh, have no Idea what this is because I'm not very far in physics, but I can provide a medal. hope you don't mind

OpenStudy (anonymous):

suppose both are intially at rest ..and moving with velocity v (both moving with equal velocity) opposite charGe attract with force F F=(q1q2/d^2)9*10^9 F=Ma a=q1q2/d^2M v=u+at v+0+(q1q2/d^2M)t t=vd^2M/Q1Q2

OpenStudy (anonymous):

So how come energy doesn't derive the same answer you got using forces?

OpenStudy (anonymous):

The kinetic energy is equal to the DIFFERENCE between the potential energy at a given point and the potential energy at the starting point.

OpenStudy (vincent-lyon.fr):

@jordanlovato It went wrong here : v(x) = sqrt( 2kq^2/mx ) Because you have forgotten the electric potential energy at the starting point, namely: kq²/d

OpenStudy (vincent-lyon.fr):

This question is quite difficult because you need the value of a generalised integral.

OpenStudy (anonymous):

Another issue is that as they meet the velocity of approach becomes infinite! You will be integrating a singularity.

OpenStudy (anonymous):

Okay, second attempt at this.. U(x) = qV(0) - qV(x) V(0) = kq^2/d V(x) = kq^2/d-2x -> both particles have moved x distance closer to each other U(x) = kq^2(2x/d(d-2x)) U(x) = KE(x) kq^2(2x/d(d-2x)) = 1/2mv(x)^2 v(x) = sqrt(2kq^2/m) * sqrt( 2x/d(d-2x) ) v(x) = dx/dt t = Int[0,1/2d] 1/v(x) dx t = sqrt(m/2kq^2) * Int [0,1/2d] sqrt(d(d-2x)/2x)dx Am I getting any closer?

OpenStudy (vincent-lyon.fr):

Keep x as the distance that will vary from d to 0. v² will be proportional to 1/x - 1/d To keep v positive , choose v = –dx/dt

OpenStudy (vincent-lyon.fr):

In the end, I think this integral is needed: http://www.wolframalpha.com/input/?i=integrate+%281%2Fx-1%29^%28-1%2F2%29+dx+from+x%3D0+to+1

OpenStudy (anonymous):

Cool, thank you sir!